OpenStudy (anonymous):

Can a right triangle be scalene? I don't think so but clarification would be much appreciated :D

OpenStudy (mathstudent55):

What does scalene trianlge mean?

OpenStudy (anonymous):

Yes for example a right triangle with sides 3 4 and 5 is scalene

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@mathstudent55 when it has no equal sides no equal angles

OpenStudy (anonymous):

Wait, so yes it CAN be or yes it CANT be? lol

OpenStudy (anonymous):

Yes it can be:P

OpenStudy (anonymous):

yes it can :)

OpenStudy (mathstudent55):

Correct. Above, you already have an example that works. Sides 3, 4, 5 Using the Pythagorean theorm, 3^2 + 4^2 = 5^2 9 + 16 = 25 25 = 25 Sides 3, 4, 5 do indeed make a right triangle and are all different, so a triangle with sides 3, 4, 5 is both right and scalene.

OpenStudy (anonymous):

Thanks. I understand now :)

OpenStudy (mathstudent55):

wlcm